### v3

# LoTLIP: Improving Language-Image Pre-training for Long Text Understanding

Abstract
Understanding long text is of great demands in practice but beyond the reach of most language-image pre-training (LIP) models. In this work, we empirically confirm that the key reason causing such an issue is that the training images are usually paired with short captions, leaving certain tokens easily overshadowed by salient tokens. Towards this problem, our initial attempt is to relabel the data with long captions, however, directly learning with which may lead to performance degradation in understanding short text (e.g., in the image classification task). Then, after incorporating corner tokens to aggregate diverse textual information, we manage to help the model catch up to its original level of short text understanding yet greatly enhance its capability of long text understanding. We further look into whether the model can continuously benefit from longer captions and notice a clear trade-off between the performance and the efficiency. Finally, we validate the effectiveness of our approach using a self-constructed large-scale dataset, which consists of $100M$ long caption oriented text-image pairs. It is noteworthy that, on the task of long-text image retrieval, we beat the competitor using long captions with 11.1% improvement (i.e., from 72.62% to 83.72%). We will release the code, the model, and the new dataset to facilitate the reproducibility and further research. The project page is available at here.
1 Introduction
Figure 1: Illustration of the impacts of long v.s. short captions on image-language pre-training, as observed in the cross-attention maps of CLIP. Training images are usually paired with short captions, leaving certain tokens (e.g., garden token) easily overshadowed by salient tokens (e.g., castle token). Fortunately, the usage of long captions can help bring the overshadowed tokens back into the light, and this phenomenon is not influenced by the order of tokens within the sentence.
Understanding long texts is a vital skill in natural language processing (NLP), extensively utilized across various domains, including text generation [22, 30]. Building on this foundation, in the multi-modality field, some text-to-image works (e.g., DALLE-3 [1] and Pixel-art [4, 5]) employ pre-trained captioners (e.g., LLaVA [20]) to generate more accurate and detailed captions—in other words, long captions—for images. These long captions significantly reduce the generative model’s tendency to hallucinate, thereby enhancing the quality of text-image alignment. However, in the realm of language-image pre-training, scant research has been conducted on modeling long texts in a way that effectively aligns with image representations. Despite the lack of exploration, this is a critical problem with practical demands because multi-modality models are increasingly common in applications that require nuanced understanding of textual descriptions corresponding to images.
There are two primary challenges hinder the effective integration of long-text understanding in language-image pre-training. The first one is the lack of large-scale long-caption image-text paired datasets. Most existing datasets focus on short captions (i.e., average text length in CC12M [3] is about 17), which limits the model’s exposure to longer text forms. Consequently, the models trained on these datasets tend to neglect certain tokens that are easily overshadowed by salient tokens. As shown in Fig. 1, the model trained on short captions can be good at understanding the content of short captions (i.e., garden and castle). But when increasing the length of caption, we can see that ‘garden token’ is overshadowed by ‘castle token’, even if we move the ‘garden token’ front. This bias towards the salient tokens (‘castle token’ in Fig. 1) of texts can severely restrict the model’s ability to comprehend and generate responses based on the full context of longer inputs. The second challenge is the token length limitation of the text encoder. While directly increasing the token number limitation that a model can process appears to be a straightforward solution for accommodating longer texts, it does not automatically translate into better understanding for long captions. The fundamental issue remains the model’s inability to effectively interpret long texts, primarily due to the lack of appropriate training data that includes long captions.
To address these challenges, we have undertaken an extensive project to re-caption 100 million data with long captions, aiming to enrich the training environment for our models. This initiative allows us to explore the effects of increased text length in image-text pre-trained models and its impacts on model performance. Based on these experiments, we empirically confirm that the key reason causing such an issue is that the training images are usually paired with short captions, leaving certain tokens easily overshadowed by salient tokens. However, directly learning with long captions may improve the long-text understanding of image-text pre-trained models, but lead to performance degradation in understanding short texts (e.g., in the image classification task) as shown in Fig. 2. After integrating corner tokens to aggregate diverse textual information, we successfully enable the model to regain its original proficiency in understanding short texts while significantly improving its ability to comprehend long texts. We also explore whether the model can continue to benefit from longer captions and observe a clear trade-off between performance and efficiency. It is noteworthy that, on the task of long-text image retrieval, we beat Long-CLIP [35], a competitor using long captions for fine-tuning, with 11.1% improvement (i.e., from 72.62% to 83.72%).
2 Related work
2.1 Language-Image Pre-training
Recently, using language-image pre-trained models to do zero-shot prediction has attracted a lot of attention. CLIP [23] and ALIGN [14] demonstrate contrastive pre-trained models can learn rich visual-language correspondence knowledge from large-scale image-text pairs on the Internet and achieve good performance on zero-shot predictions, including image-text retrieval [18] and classification [9]. Following their success, various studies [31, 17, 16] have been devoted to improving image-text alignment. Among them, FILIP [31] focuses on fine-grained expressiveness between text tokens and image patches by modifying the training loss. While LiT [33] finds that apply contrastive-tuning to the pre-trained models with locked image encoder and unlocked text encoder can further improve the alignment. It is recognized that larger batch size brings better performance. For this reason, SigLIP [34] proposes to replace the softmax normalization among the standard contrastive loss with the sigmoid loss to scale up training batch size. In addition, LaCLIP [11], RLEG [36] and some other works [28, 13, 21] utilize multi-modality generative models to improve data quality for enhancing pre-training.
2.2 Long-text Understanding
Detailed long texts are necessary for many artificial intelligence tasks (e.g., human-computer interaction). Therefore, modeling and parsing long texts has become one of the most important research directions in natural language processing. Many studies in fields such as text generation [22, 30] have shown that advanced transformer structures have the ability to interpret long texts. However, in the field of language-image pre-training, research on using long-text descriptions of images to enhance multimodal representations is still very scarce. DreamLIP [37] utilizes multi-modality large language model to re-caption image data with detailed descriptions and then use them in language-image pre-training. In fact, during training, DreamLIP randomly extracts sub-captions from the detailed description for training and does not completely use all the information of the long text. One of the reasons why previous language-image pre-training rarely directly applied long texts for training is the text encoder of traditional CLIP [23] is restricted by the token number limit ($\leq 77$). Then, Long-CLIP [35] firstly introduces long captions into CLIP model to finetune, where the model is pre-trained on short-text-image datasets. The fine-tuning process equips the model with the ability to comprehend long texts, albeit at the expense of its proficiency in understanding shorter texts. In contrast, we incorporate long captions during the pre-training stage, which can not only enable the model to regain its original competency in interpreting short texts but also significantly improves its understanding of long texts. We further explored the trade-off between the benefit of long texts to the model and the training efficiency.
3 Preliminary of Language-Image Pre-training
Language-image pre-training models, i.e., CLIP [23] and LiT [33], typically consist of two parts: an image encoder and a text encoder. In the training stage, the language-image model takes image-text pairs as input, and the image encoder and text encoder are used to extract embeddings from images and texts, respectively. Then, two encoders are trained with contrastive objectives, ensuring that paired image and text embeddings are close in the embedding space, while unpaired pairs are far apart. Specifically, in the pre-training stage, a batch of images $\{I_{1},I_{2},\cdots,I_{N}\}$ and the corresponding short texts $\{T_{1},T_{2},\cdots,T_{N}\}$ are randomly sampled from the pre-training dataset. Each image and its corresponding text are treated as a positive pair, while others in the same batch are negative pairs. Then, the image encoder extracts image global feature $\mathbf{v}^{i}$ from the $i$-th image $I_{i}$ within the batch, while the text encoder obtains text feature $\mathbf{t}^{j}$ from the $j$-th text $T_{j}$. These two encoders are then optimized using contrastive loss $\mathcal{L}$, which consists of image-to-text loss $\mathcal{L}^{i2t}$ and text-to-image loss $\mathcal{L}^{t2i}$. It can be formulated as follows:
$\mathcal{L}=\mathcal{L}^{i2t}+\mathcal{L}^{t2i},$ (1)
$\mathcal{L}^{i2t}=-\sum_{i=1}^{N}\log\frac{\exp\left(\cos\langle\mathbf{v}^{i},{\mathbf{t}^{i}}\rangle/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\langle\mathbf{v}^{i},{\mathbf{t}^{j}}\rangle/\tau\right)},$ (2)
$\mathcal{L}^{t2i}=-\sum_{i=1}^{N}\log\frac{\exp\left(\cos\langle\mathbf{t}^{i},{\mathbf{v}^{i}}\rangle/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\langle\mathbf{t}^{i},\mathbf{v}^{j}\rangle/\tau\right)},$ (3)
where $\tau$ is a learnable temperature parameter, and $\cos\langle\cdot,\cdot\rangle$ means the cosine similarity between two normalized feature vectors.
4 Long Texts in Language-Image Pre-training
Currently available image-text pair datasets, e.g.CC12M [26], typically include short texts that have an average length of approximately 17 tokens. Language-image models pre-trained on these datasets perform well on short-text comprehension tasks. However, we find that they struggle to comprehend long texts for text-image alignment. Concretely, they tend to overlook or neglect some tokens or sub-captions in long texts, as shown in Fig. 1. A potential reason for such a situation is the lack of long-text-image pairs in pre-training. Thus, we collected and re-captioned 100M images with long texts. In this section, we provide details of re-captioning and explore the usage of long texts in language-image pre-training.
4.1 Long Text-Image Pair Dataset
Table 1: Dataset details of long-text-image retrieval and short-text-image retrieval tasks. We use BERT tokenizer for tokenization. ShareGPT4V-1k and 10k are selected from the ShareGPT4V dataset. For DCI and IIW, all images with human-authored long descriptions are used while evaluating.
[…]
Long-text-image Retrieval Dataset
DCI [29] 7,805 7,805 10.81 172.73
IIW [12] 612 612 10.16239.73
ShareGPT4v-1k [6] 1,000 1,000 8.15 173.24
ShareGPT4v-10k [6] 10,00010,000 8.24 173.66
Short-text-image Retrieval Dataset
MSCOCO [18] 5,000 25,000 1.0 11.77
Flickr30k [32] 1,000 5,000 1.0 14.03
Training Dataset.
To construct long text-image pairs for language-image pre-training, we re-captioned 100 million images with long texts. Specifically, we collected the images from CC3M [26], CC12M [26], YFCC15M [27], LAION [25], and COYO [2] dataset. Then, we instructed three multi-modality large language models (MLLMs), i.e., InstructBLIP [7], LLaVA [19], and ShareGPT4V [6] to generate diverse and descriptive long texts based on the collected images. In this step, we used “Describe the image in detail.” as the text prompt, following DreamLIP [37]. Finally, each image in the collected datasets is paired with four texts: a raw text from the original dataset and three re-captioned long texts. The raw texts exhibit an average length of approximately 18 tokens, whereas the re-captioned long texts consist of around 136 tokens.
Evaluation Dataset.
Most zero-shot evaluation tasks for language-image pre-training primarily rely on short textual input. For example, in short-text-image-retrieval tasks, the textual inputs contain fewer than 15 tokens on average, as shown in Table 1. Given the short texts, these tasks are not suitable to access the long text comprehension ability of pre-trained models. Therefore, we collected long text-image pairs from DCI [29], IIW [12], and ShareGPT4V [6] datasets for constructing long-text-image retrieval evaluation task. Specifically, for DCI and IIW datasets, we use images with human-annotated long descriptions for retrieval. For ShareGPT4V dataset, we sample 1,000 and 10,000 data from ShareGPT4V dataset to construct ShareGPT4V-1k and ShareGPT4V-10k retrieval dataset, respectively. The ShareGPT4V-1k dataset is constructed following Long-CLIP [35]. All images in ShareGPT4V-1k and ShareGPT4V-10k are from SA-1B [15] dataset, and all long texts are generated by ShareGPT4V-Captioner [6]. As shown in Table 1, each image within these datasets has one paired long text, which consists of more than 8 sub-captions and 170 tokens on average. Here, a sub-caption is a complete sentence ending with a period. The long-text-image retrieval task necessitates the alignment of text features from long texts with the image features from the corresponding images, which thereby evaluates the model’s ability to comprehend long texts.
Figure 2: The influence of text length. A significant improvement is observed across all tasks when we added one randomly sampled sub-caption from generated texts to the pre-training stage. As the number of sub-captions increases, the performance of the pre-trained model on long-text-image retrieval tasks consistently improves and becomes stable (a). However, there is a performance degradation in MSCOCO retrieval task (b) and ImageNet classification task (c).
4.2 Exploring the Influence of Text length
Long-CLIP [35] enables long text processing ability by fine-tuning CLIP model with long text-image pairs. However, the benefits and drawbacks of using long texts in pre-training are still unknown. It is also unknown how the length of the text affects the performance of the pre-trained model. To explore the influence of using texts in different lengths for pre-training, we change the length of long texts by selecting different numbers of consecutive sub-captions. Each sub-caption in the long texts has an average of approximately 22 tokens. The experiment is conducted on 3M scale dataset and the results are illustrated in  Fig. 2. Sub-caption number equals 0 indicates that the model is trained without long texts. When introducing one sub-caption, a noticeable gain is achieved across all tasks. The results also indicate that longer texts, composed of more sub-captions, enhance the performance of the pre-trained model on long-text-image retrieval tasks. But it tends to stabilize when the number of sub-captions reaches three. It confirms that using long text-image pairs for pre-training improves pre-trained models’ understanding of long texts. However, longer texts in language-image pre-training degrades the performance on short-text-image retrieval and image classification tasks.
4.3 Method
[…]
As presented in Sec. 4.2, directly incorporating long texts into pre-training negatively impacts the performance of model in short-text comprehension tasks. In order to find a solution that well balances both long and short texts, we design to add extra text tokens for text encoders, termed corner tokens, which can aggregate diverse text features. This strategy benefits the class (i.e., [CLS]) token by extracting more representative features for long and short text. Next, we will describe the details.
Corner Tokens.
[…]
$\mathcal{A}(q,k)=\left\{\begin{array}[]{ll}0,&\text{if }(k\in\mathcal{C}\text{ or }q,k\in\mathcal{C}\cup\text{\{[{CLS}]\}})\text{ and }q\neq k\\ 1,&\text{otherwise}\end{array}\right.$
[…]
Optimization.
The short-text-image contrastive loss $\mathcal{L}_{\text{short}}$ is calculated in the same way as Eq. 1. Meanwhile, the long-text-image contrastive loss $\mathcal{L}_{\text{long}}$ between image global feature $v$ and $\mathbf{t}_{g},\mathbf{t}_{c_{1}},\mathbf{t}_{c_{2}},...,\mathbf{t}_{c_{m}}$ of long text is as follows:
$\mathcal{L}_{\text{long}}=\mathcal{L}_{\text{long}}^{i2t}+\mathcal{L}_{\text{long}}^{t2i},$ (4)
$\mathcal{L}_{\text{long}}^{i2t}=-\sum_{i=1}^{N}(\log\frac{\exp\left(\cos\langle\mathbf{v}^{i},{\mathbf{t}_{g}^{i}}\rangle/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\langle\mathbf{v}^{i},\mathbf{t}_{g}^{j}\rangle/\tau\right)}+\sum_{k=1}^{m}\log\frac{\exp\left(\cos\langle\mathbf{v}^{i},\mathbf{t}_{c_{k}}^{i}\rangle/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\langle\mathbf{v}^{i},\mathbf{t}_{c_{k}}^{j}\rangle/\tau\right)}),$ (5)
$\mathcal{L}_{\text{long}}^{t2i}=-\sum_{i=1}^{N}(\log\frac{\exp\left(\cos\langle\mathbf{t}_{g}^{i},\mathbf{v}^{i}\rangle/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\langle\mathbf{t}_{g}^{i},\mathbf{v}^{j}\rangle/\tau\right)}+\sum_{k=1}^{m}\log\frac{\exp\left(\cos\langle\mathbf{t}_{c_{k}}^{i},\mathbf{v}^{i}\rangle/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\langle\mathbf{t}_{c_{k}}^{i},\mathbf{v}^{j}\rangle/\tau\right)}),$ (6)
The total training loss is $\mathcal{L}_{\texttt{LoTLIP}}=\mathcal{L}_{\text{long}}+\mathcal{L}_{\text{short}}$.
5 Experiments
5.1 Implementation Details and Datasets
Pre-training Datasets. As presented in Sec. 4.1, we collected 100M data from five publicly available image-text pair datasets and re-captioned the collected images with long texts. Based on this dataset, we construct 4 scales of pre-training data: (1) 3M, including CC3M. (2) 12M, including CC12M. (3) 30M, including CC3M, CC12M, and YFCC15M. (4) 100M, including all re-captioned data. We conduct ablation studies to validate our model on the 3M scale pre-training data. The performance of LoTLIP pre-trained with 12M and 30M scale datasets is shown in the Supplementary Material.
Downstream Datasets. To assess the ability of the pre-trained models on short-text and long-text understanding, we select 3 downstream tasks for evaluation under the zero-shot setting, including long-text-image retrieval, short-image-text retrieval, and image classification. For long-text-image retrieval, we present the Recall at 1 (R@1) metric of the pre-trained models on DCI, IIW, and ShareGPT4V-1k, and ShareGPT4V-10k long-text-image retrieval tasks. For short-image-text retrieval, we evaluate on MSCOCO [18] and Flickr30k Caption [32] and report Recall at 1/5 (R@1/5) metric for comparison. For image classification, we use ImageNet1k [9] for evaluation and present top-1 accuracy (Acc@1) on image classification. Following [23], we use class names incorporated with pre-defined text prompts as text inputs for zero-shot classification.
Implementation Details. Following LiT [33], we use a vision transformer pre-trained on ImageNet 21K as the image encoder and Bert [10] as the text encoder. The architecture of the image encoder is ViT-B/16. We report other variants of vision transformers in the Supplementary Material. The images are resized to 224$\times$ 224. The maximum text token length is set to 128 unless specifically stated. Three consecutive sub-captions are randomly selected to from long texts as text input. We train 10 epochs on the 3M and 100M scale datasets. For the 3M dataset, the batch size is set to 2560, while that of 100M is set to 16384. The other pre-training hyperparameters are under the same setting, e.g.learning rate, warmup steps, and weight decay.
5.2 Ablation Studies
Exploring the Influence of Token Number Limitation
Figure 4: Influence of token number limitation on LoTLIP. The performance of the pre-trained model on different tasks improves when the token number limitation increases up to 192, which exceeds the commonly used 77. Meanwhile, the FLOPs of the text encoder (red stars) rapidly increase with the text token number limitation.
Token number limitation is the maximum length of text input that the text encoder can process at a time. For language-image pre-training models, the token number limitation of text encoder can be arbitrarily set to any positive integer, which is typically set to 77 [23, 34, 37]. The benefits derived from using long texts in pre-training might be constrained by the 77 token limitation. To investigate this hypothesis, we incrementally raise the token number limitation from 32 to 512. The results are illustrated in Fig. 4, which indicate a limitation of 77 tokens is insufficient for a model training with long texts. On DCI and ShareGPT4V-10k retrieval tasks, the best results are observed when the token number limitation is set to 192. Meanwhile, the highest performance on the IIW retrieval task is reached when the text token length limitation is 256. Since the average text length of the evaluation datasets is less than 240 tokens, as shown in Table 1. When the text token length limitation gets larger than 256, these datasets can’t well prove the long-text understanding ability of the model. In Fig. 4, we also illustrate the FLOPs of the text encoder, which increases with the text token number limitation. To balance the training efficiency and performance, we set the token number limitation to 128 for the text encoder of LoTLIP.
Table 2: Analyze the effectiveness of LoTLIP in language-image pre-training with long texts. The architecture of the image encoder is ViT-B/16. T2I and I2T indicate R@1 on text and image retrieval, respectively. We use 3M scale dataset for pre-training. “✓" indicates we add long texts in the training stage.
[…]
DCI IIW ShareGPT4v-10kMSCOCO ImageNet
T2I I2T T2I I2T T2I I2T T2I I2T Acc@1
LiT - 27.14 24.13 65.20 58.50 32.73 27.01 24.0734.20 43.76
LiT ✓ 47.96 44.92 84.97 81.70 73.66 66.73 30.06 43.52 48.87
LiT+Long-CLIP* ✓ 47.21 46.61 83.82 83.01 74.60 67.49 26.74 43.78 46.82
LoTLIP49.4647.8284.9783.3376.4969.7231.5946.5650.34
    * We apply the primary Component matching strategy proposed by Long-CLIP [35] to LiT and train the model with the same training setting using 3M data for fair comparison.
Compare LoTLIP with other methods in pre-training with long texts
LoTLIP aims to enhance the long-text and short-text comprehension ability of language-image pre-training. To demonstrate the effectiveness of LoTLIP, we compare it with other approaches that train with long text-image pairs. The first approach is directly using long texts in the training stage of LiT. The second approach is based on another contrastive learning method related to long texts, namely Long-CLIP [35]. Concretely, we incorporate the primary component matching strategy and losses proposed by Long-CLIP to LiT (LiT+Long-CLIP). For a fair comparison, all models are trained with 3M scale dataset. As shown in Table 2, directly using long texts in pre-training procedural significantly improves LiT over all tasks. Moreover, LiT+Long-CLIP exceeds LiT on the long-text-image retrieval tasks when both models incorporate long texts during pre-training. But on short-text comprehension tasks, i.e., short-text-image retrieval and image classification, the performance of LiT+Long-CLIP is inferior to that of LiT. Instead, LoTLIP improves LiT and LiT+Long-CLIP on both long and short text comprehension tasks. Concretely, LoTLIP surpasses LiT and LiT+Long-CLIP by 1.97% and 1.51% on average over three long-text-image retrieval tasks. Besides, LoTLIP improves LiT by 2.29% and 1.47% on MSCOCO retrieval and ImageNet classification, respectively. The results demonstrate that LoTLIP benefits from long texts and corner tokens, thereby exhibiting a better understanding of the long and short texts.
Table 3: Analyze the influence of the number of corner tokens and the attention mask mechanism. We use 3M scale dataset for training. The architecture of the image encoder is ViT-B/16.
[…]
Mask Long-text-image Retrieval Short-text-image Retrieval Classification
DCI IIW ShareGPT4v-10k MSCOCO ImageNet
T2I I2T T2I I2T T2I I2T T2I I2T Acc
0 - 47.96 44.92 84.97 81.7073.66 66.73 30.06 43.52 48.87
1 ✓ 49.57 46.55 84.97 82.6874.91 68.41 31.51 45.68 49.62
2 ✓ 49.46 47.82 84.97 83.33 76.49 69.7231.5946.56 50.34
3 ✓ 49.58 47.70 87.0984.3176.5170.20 31.60 46.48 50.36
4 ✓ 49.5848.30 86.76 84.1576.25 70.1431.88 47.7050.59
2 - 48.61 47.17 86.11 81.8676.14 69.31 31.34 47.70 49.88
2 ✓ 49.46 47.82 84.97 83.33 76.49 69.7231.5946.56 50.34
Table 4: Zero-shot evaluation of different models on long-text-image retrieval tasks. T2I and I2T indicate R@1 on text and image retrieval, respectively.
[…]
Short Long T2I I2T T2I I2T T2I I2T T2I I2T
3M - FILIP [31] 10.85 11.36 31.54 29.0826.50 26.80 8.94 8.64 19.21
3M - LaCLIP [11] 14.84 14.71 41.01 38.89 40.90 37.10 15.8114.84 27.26
3M - SigLIP [34] 11.66 13.11 29.25 29.5827.30 25.10 9.92 9.30 19.40
3M - LiT [33] 27.14 24.13 65.20 58.50 63.60 56.80 32.73 27.01 44.38
3M LoTLIP 49.46 47.82 84.97 83.33 93.20 90.00 76.49 69.72 74.37
400M - CLIP [23] 45.45 43.01 88.24 87.58 84.5079.80 60.2256.16 68.12
100M - LiT [33] 41.78 40.90 88.07 82.68 86.00 80.00 61.41 50.69 66.44
700M -ALIGN [14]56.5457.41 92.65 90.68 86.30 85.30 65.13 62.73 74.59
12B - SigLIP [34]57.78 56.22 91.99 91.01 85.80 83.40 83.4063.08 76.59
400M 1M Long-CLIP* [35] 47.4344.1889.2286.93 90.60 87.40 73.16 62.03 72.62
100M
LoTLIP 62.10 61.06 93.95 92.48 96.50 95.50 86.84 81.40 83.72
    * Long-CLIP fine-tunes pre-trained CLIP model with ShareGPT4v-10k data from ShareGPT4v datasets except ShareGPT4v-1k.
Implementation of Corner Tokens
In this part, we study the influence of the attention mask mechanism and the number of corner tokens on different downstream tasks as shown in Table 3. LoTLIP pre-trained without the pre-defined attention mask encounters performance degradation on most tasks compared to when using the pre-defined attention mask. It indicates that direct interaction between the CLS token and corner tokens limits their ability to aggregate diverse textual features. On short-text-image retrieval and image classification tasks, the performance of LoTLIP improves when introducing more corner tokens. It proves that the corner tokens help with enhancing the short text understanding ability of LoTLIP. When the number of corner tokens is set to more than two, the performance improvement across all tasks is relatively small. Thus, we use two corner tokens in LoTLIP.
5.3 Main Results
We compare LoTLIP with the state-of-the-art methods on downstream tasks involving long texts and short texts. The experimental results are shown in Table 4 and Table 5, respectively. On 3M data scale, LoTLIP significantly improves the state-of-the-art methods on all tasks. Specifically, LoTLIP improves the second best method LiT by 29.99% on average over four long-text-image retrieval tasks. Moreover, LoTLIP improves the second competitor LiT by 6.58% and 10.98% on image classification task and short-text-image retrieval tasks, respectively. It proves that LoTLIP significantly enhances language-image model for understanding short and long captions by involving long captions in pre-training and incorporating corner tokens in text inputs. It is worth noting that LoTLIP trained with 100M data exceeds all state-of-the-art methods on long-text-image retrieval tasks, even though these methods are pre-trained on a larger scale of data. Concretely, LoTLIP (trained on 100M data) exceeds SigLIP (trained with 12B) by an average of 7.13% over four long-text-image retrieval tasks. Moreover, compared to Long-CLIP, which first uses long texts in CLIP for fine-tuning, LoTLIP improves averaged performance by 11.1% on long-text-image retrieval tasks.
Table 5: Zero-shot evaluation of different models on short-text-image retrieval and classification tasks.
[…]
Short Long Acc. R@1R@5 R@1R@5 R@1R@5 R@1R@5
3M - FILIP [31] 18.69 14.98 34.88 11.86 28.98 30.40 56.80 20.76 44.20
3M - LaCLIP [11] 21.50 18.94 40.40 12.42 31.04 37.00 63.90 29.32 56.04
3M - SigLIP [34] 21.28 16.30 37.36 13.22 31.65 31.00 61.10 23.7648.64
3M - LiT [33] 43.76 34.20 61.52 24.07 48.37 61.30 89.50 48.06 75.36
3M LoTLIP 50.34 46.56 72.02 31.59 57.65 75.20 94.20 58.20 83.58
400M - CLIP [23] 68.34 51.68 76.72 32.70 57.76 82.20 96.60 62.14 85.72
100M - LiT [33] 73.70 51.92 75.72 32.74 57.84 80.00 95.90 60.86 84.62
700M - ALIGN [14] 65.89 60.42 82.50 42.36 67.38 88.90 98.00 74.12 92.40
12B - SigLIP [34] 76.04 65.46 86.22 47.14 72.10 89.10 98.00 74.66 92.30
400M 1M Long-CLIP [35] 67.10 57.28 80.78 40.34 65.92 85.90 98.50 70.66 90.60
100M
LoTLIP 72.1659.66 81.50 38.06 63.81 86.90 97.80 65.22 87.98
6 Conclusion
In this work, we empirically confirm that a key issue arises because training images are typically paired with short captions, which can cause certain tokens to be overshadowed by more salient ones. To address this issue, our initial strategy involved relabeling the data with long captions. However, directly learning from these long captions might lead to degraded performance in tasks requiring an understanding of short text, such as image classification. Subsequently, by incorporating corner tokens to aggregate diverse textual information, we are able to help the model regain its original proficiency in understanding short texts while significantly enhancing its capability to comprehend long texts. We also investigated whether the model could continue to benefit from longer captions and observed a clear trade-off between performance and efficiency. Finally, we validated the effectiveness of our approach using a self-constructed large-scale dataset, which consists of 100 million long-caption-oriented text-image pairs. It is noteworthy that in the task of long-text image retrieval, our method outperforms the long caption-related competitor, Long-CLIP, with a significant improvement of 11.1%. We will release the code, the model, and the new dataset to facilitate the reproducibility and further research.